Please tell me if I am right and if not, show me the answer please. Find the length of the following segment created by these pairs of endpoints!
(1,2) and (7,2) I believe it goes like this (1+7)/2=8 and (2+2)/2=4/2=2 Solution is (8,2) Correct?

Answer by Alan3354(69443)

Please tell me if I am right and if not, show me the answer please. Find the length of the following segment created by these pairs of endpoints!
(1,2) and (7,2) I believe it goes like this (1+7)/2=8 and (2+2)/2=4/2=2 Solution is (8,2) Correct?
----------------------------
You found a point, not the length.
Since y is 2 for both given points, the length is 6 (7-1).
------------
It appears you were finding the midpoint, but you didn't divide the x values by 2. (1+7)/2 = 4, not 8.